OpenStudy (anonymous):

check my work to see if i did this right? a + 5 = -5a + 5 Add -5a to both a's -4a + 5 = 5 Subtract 5 from both 5's -4a = 0 Divide -4a to both sides a = 0 i think did i do this right? please help me. i really want to know how to do these. i will give a medal for the best response... :)

OpenStudy (anonymous):

Yes this is correct. -4a divided by 0 is always zero.

OpenStudy (anonymous):

yay! thank you so much!! could you help me with this one, also? i promise it will be the last one... p - 1 = 5p + 3p - 8 p - 1 = 8p - 8 Subtract p to both sides -1 = 7p - 8 Add 8 to both sides -7 = 7p Divide 7p to both sides p = -1 im sorry. i do online school and algebra is really hard for me, and now we are learning the two side equations. i just have a lot of trouble with it.